Qd= 25000-2p


Qs= 10000-1p


Calculate the market equlibrium level of price and quantity for housing unit?


1
Expert's answer
2021-05-06T15:59:51-0400

The market level of equilibrium price and quantity is determined by putting demand equation equal to supply equation. This is because equilibrium is attained where demand curve intersects supply curve.

"Q_d=25000-2P"

"Q_s=10000-1P"


Where,

"Q_d=quantity\space demanded"

"Q_s=quantity\space supplied"

"P=price"


"putting\space Q_d=Q_s"

"25000-2P=10000-1P"

"25000-10000=2P-1P"

"P=15000"

The equilibrium price is 15000


Now, putting P=15000 in the demand and supply equation gives,

"Q_d=25000-2P"

"Q_d=25000-2(15000)"

"Q_d=-5000"


"Q_s=10000-1P"

"Q_s=10000-15000"

"Q_s=-5000"


The equilibrium quantity is -5000.

 

The equilibrium quantity is negative because the supply equation also shows an inverse relationship between price of housing units and quantity supplied which is different from the general case of inverse relationship between price of a good and its quantity supplied.